What is the perimeter and area of a rectangle with a side of 7ft and a diagonal of 25ft?
Answer by nyc_function(260)

We are given one side and a diagonal.
We can think of the side as one of the legs of a right triangle and the diagonal as the hypotenuse of the same right triangle.
We need to find the other leg as step one.
To do this, we use the Pythagoren Theorem from geometry.
(leg)^2 + (leg)^2 = (hypotenuse)^2
(leg)^2 + 7^2 = 25^2
(leg)^2 + 49 = 625
(leg)^2 = 625 - 49
(leg)^2 = 576
Take square root of both sides.
leg = 24
So, the other side of the rectangle is 24 feet.
We now use two formulas:
To find the area, use A = length times width.
To find the perimeter use, P = 2(length) + 2(width).
A = 7 times 24
A = 168ft^2
===================================================
P = 2(24) + 2(7)
P = 48 + 14
P = 62 feet